OpenStudy (anonymous):

LaGrange Multipliers: I figured out how to do these a while back but now I can't remember and this book is horrible. Just a basic problem: f(x,y) = x+3y+2, x^2+y^2=10 Find the max and min.

OpenStudy (anonymous):

Is the min (-1,-3) and the max (1,3)?

OpenStudy (tkhunny):

Create \[f(x,y,\lambda) = x + 3y + 2 -\lambda(x^{2}+y^{2}-10)\] Find \[\frac{\partial f}{\partial x}\]\[\frac{\partial f}{\partial y}\]\[\frac{\partial f}{\partial \lambda}\] Then what?

OpenStudy (anonymous):

Nevermind.

OpenStudy (tkhunny):

Too late. It was already fun!

OpenStudy (anonymous):

Braha. These are fun once I remember that it is -c not +c.